OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!! Finish these three solutions to this equation. x-y=5 (__, 0), (__, 3), (__, 5)

OpenStudy (mertsj):

Add y to both sides of the equation and you will see that x is 5 more than y

OpenStudy (mertsj):

So just add 5 to all the y values to get the x values

OpenStudy (anonymous):

1) x-0=5 2) x-3=5 x=5+3 3) x-5=5 x=5+5 (I tried my best not to give the answer. hope this would help)